# Large Language Models Meet Knowledge Graphs for Question Answering: Synthesis and Opportunities

### v1

@@ 3.1.3 Aligning with Complex QA
In the approaches of synthesizing LLMs and KGs for QA where KG plays the role of background knowledge, the retrieved relevant knowledge from the factual knowledge graph can reconcile knowledge conflicts from the multiple documents for multiple-doc QA (Wang et al., 2024d). The cross-modal reasoning can facilitate the cross-modal interaction and alignment for multi-modal QA (Suri et al., 2024). Additionally, the question decomposition of multi-hop QA can be augmented by fusing the knowledge from LLMs and KGs, which further facilitates iterative reasoning for generating the accurate final question (Saleh et al., 2024; Cao et al., 2024). The RAG (Roy et al., 2024) and KG-RAG (Sanmartin, 2024) can also improve the capabilities of LLMs in understanding the user’s interactions for generating accurate answers for multi-run and conversational QA. However, the key technical challenge behind this methodology is how to retrieve the relevant knowledge from large-scale KGs and then effectively fuse with LLMs without inducing knowledge conflicts?
@@ 3.2.1 Offline KG Guidelines
In offline KG guidelines, KG supplies potential paths or subgraphs before the reasoning process by the LLM, from which the LLM selects the most relevant path for reasoning. For instance, SR (Zhang et al., 2022) trains a subgraph retriever that operates independently from the downstream reasoning process, while Keqing (Wang et al., 2023) decomposes complex questions using predefined templates, retrieves candidate entities and triples from a KG. EtD (Liu et al., 2024a) initially uses a GNN to identify promising candidates and extracts relevant fine-grained knowledge pertinent to the questions, and then creates a knowledge-enhanced multiple-choice prompt to guide LLM for generating the final answer. Recent studies have been exploring the application of novel formats of guidelines. GCR (Luo et al., 2024a) first converts a KG into a KG-Trie and then develops a graph-constrained decoding and a lightweight LLM to generate multiple reasoning paths and candidate answers. KELDaR (Li et al., 2024b) introduces the question decomposition and atomic retrieval modules to extract implicit information and retrieves the relevant subgraphs from KG to augment LLM for QA.
@@ 3.2.2 Online KG Guidelines
This paradigm emphasizes that the guidance of the KG directly involves the LLM’s reasoning process, rather than merely relying on existing static knowledge. For example, Oreo (Hu et al., 2022) uses a contextualized random walk across a KG and conducts a single step of reasoning through the specific layers. KBIGER (Du et al., 2022) considers the ($k$-1)-th reasoning graph to construct the $k$-th instruction for reasoning and corrects erroneous predictions of intermediate entities. LLM-ARK (Huang, 2023) approaches reasoning tasks as sequential decision-making processes and employs Proximal Policy Optimization (PPO) for optimization. ToG (Sun et al., 2024a) allows LLMs to iteratively perform beam search over KGs, by which the most promising reasoning paths and the most likely reasoning outcomes are generated. In contrast, ToG-2 (Ma et al., 2025b) utilizes entities as intermediaries to guide LLMs toward precise answers based on iterative retrieval between documents and KGs. KG-CoT (Zhao et al., 2024b) leverages external KGs to generate reasoning paths for joint reasoning of LLMs and KGs to enhance the reasoning capabilities of LLMs for QA.
@@ 3.2.3 Agent-based KG Guidelines
This integration allows the Agent to leverage structured knowledge for augmenting the decision-making and problem-solving capabilities of LLMs. KG-Agent (Jiang et al., 2024) is a multifunctional toolbox integrating LLMs with a KG-based executor and a knowledge memory system that autonomously selects tools and updates the memory to enhance the reasoning of LLMs over KGs. ODA (Sun et al., 2024b) incorporates KG reasoning capabilities through a global observation approach, which improves reasoning abilities by employing a cyclical paradigm of observation, action, and reflection. GREASELM (Zhang et al., 2021) effectively integrates encoded representations from LMs and GNNs by introducing several modality interaction layers to seamlessly blend structured knowledge with language contexts.
@@ 3.2.4 Aligning with Complex QA
The approaches incorporating the KGs with LLMs can enable multi-hop and iterative reasoning over the factual graph of KGs and further augment the reasoning capability of LLMs for complex QA. The challenges of knowledge retrieval, knowledge conflicts, interactions across modalities, and complex question decomposition in multi-document QA, multi-modal QA, and multi-hop QA can be easily mitigated based on joint reasoning over the factual KGs and LLMs (Suri et al., 2024; Liang et al., 2025; Qiao et al., 2024). Moreover, the joint reasoning over the factual KGs and LLMs provides the logical inference chains and anchor for LLMs to generate explainable answers with clear evidence from factual KGs (Zhao et al., 2024b). Although the joint reasoning over the factual KGs and LLMs can address the challenges of complex QA well, joint KG–LLM reasoning remains inefficient because large-scale graph traversal is computationally intensive and time-consuming. Moreover, the reasoning capabilities of KGs mainly depend on the completeness and knowledge coverage of KGs, where the incomplete, inconsistent, and outdated knowledge from KGs might induce noise or conflicts. The main challenge lies in how to improve the reasoning efficiency over the large-scale graph and reasoning capabilities under incomplete KG?
@@ 3.3.1 KG-Driven Filtering and Validation
KGs enhance the accuracy and reliability of LLM outputs by filtering and validating candidate answers with structured and verified information. For instance, ACT-Selection (Salnikov et al., 2023) filters and re-ranks answer candidates based on their types extracted from Wikidata. Q-KGR (Zhang et al., 2024c) improves the reasoning capabilities of LLMs by filtering out the irrelevant knowledge based on the ranking of the relevance score between the question and knowledge. KGs can improve the factual accuracy of generated answers, as demonstrated by KG-Rank (Yang et al., 2024), which integrates medical KGs with re-ranking techniques to enhance the credibility of generated responses. Moreover, KGR (Guan et al., 2024) autonomously extracts and validates factual statements, significantly boosting performance on factual QA.
@@ 3.3.2 KG-Augmented Output Refinement
Integrating KGs with LLMs is essential to refining the outputs of LLMs for greater clarity and accuracy. EFSUM (Ko et al., 2024) employs LLM as a fact summarizer to generate relevant summaries from KGs, thereby improving performance in zero-shot QA. InteractiveKBQA (Xiong et al., 2024) enables iterative knowledge interactions, allowing LLMs to generate logical forms and refine outputs based on user feedback. LPKG (Wang et al., 2024b) fine-tunes LLMs with KG-derived planning data to enhance the planning capabilities of LLMs for sophisticated reasoning in complex QA.
@@ 3.3.3 Aligning with Complex QA
The approaches that leverage the retrieved factual evidence from KGs for refinement and validation are designed to augment the capability of LLMs in understanding user interactions and verifying the intermediate reasoning for multi-hop QA (Chen et al., 2024a) and multi-run and conversational QA (Xiong et al., 2024), respectively. However, the common case is that the knowledge from KGs is not available or not enough to verify the intermediate results, since the factual knowledge from KGs might be incomplete (Zhou et al., 2025). In particular, the knowledge conflicts between the intermediate answer and KG facts might induce irrelevant results due to the poorly verified intermediate results. Meanwhile, the refinement and validation of results largely depend on the correctness, timeliness, and completeness of factual knowledge in KGs. Thereby, the main challenges behind this methodology are how to handle the knowledge conflicts between the intermediate answer and KG facts and incrementally update the KGs to ensure factual knowledge in KGs is up-to-date and correct?
@@ 3.4.1 Hybrid Method
In addition to acting as a single role, the approaches (comparison in Table 6) where KGs serve multiple roles when synthesizing with LLMs for QA have been investigated. For instance, KG-RAG (Sanmartin, 2024) introduces the Chain-of-Explorations (CoE) to rank and obtain the top-$k$ nodes or relationships based on vector similarity, by which the most relevant knowledge is selected and then fed into LLMs for the final answer. LongRAG (Zhao et al., 2024a) retrieves the top-$k$ most relevant chunks based on a hybrid retriever and analyzes their relevance with the query by introducing a CoT-guided filter. Furthermore, the KG-augmented prompting is introduced to LLMs for augmenting the generation of the final answer. In KG-Rank (Yang et al., 2024), multiple ranking methodsare introduced to refine the retrieved triples for augmenting the reasoning with the most relevant knowledge. FRAG (Zhao, 2024) introduces reasoning-aware and flexible-retrieval modules to retrieve reasoning paths from KGs, thereby guiding and augmenting LLM for efficient reasoning and answer generation. KGQA (Ji et al., 2024) combines the CoT-based prompting with graph retrieval to improve the retrieval quality and the multi-hop reasoning capability of LLMs for QA.
@@ 3.4.2 Optimization
To mitigate the low efficiency and high computing costs of existing methods in unifying LLMs and KGs for complex QA, several optimization techniques (extension in Appendix § D) are proposed to improve the efficiency of synthesizing LLMs and KGs. (1) Index-based optimization. It aims to accelerate the process of learning embeddings and vector storage for given questions and knowledge context. For instance, PG-RAG (Liang et al., 2024b) proposes dynamic and adaptable knowledge retrieval indexes based on LLMs that can effectively handle complex queries and improve the overall performance of RAG systems in QA tasks. (2) Prompt-based optimization. It mainly enhances the quality of the prompts based on prompt engineering. For example, KGP (Wang et al., 2024d) proposes a KG prompting approach to enhance the prompt for LLMs and optimize the knowledge retrievalby introducing the KG construction module and LLM-based graph traversal agent. (3) Cost-based optimization. It aims to minimize computation costs by reducing the number of calls to LLMs and accelerating knowledge retrieval. In particular, SPOKE KG-RAG (Soman et al., 2024) proposes a token-based optimized KG-RAG framework that integrates explicit and implicit knowledge from KG with LLM to enhance LLMs for cost-effective QA.
@@ 3.4.3 Aligning with Complex QA
The hybrid method that involves the multiple roles of KG when synthesizing LLMs and KGs can address the limitations and challenges of complex QA well. The combination of knowledge fusion, RAG, CoT-based reasoning, and ranking-based refinement is capable of accelerating the complex question decomposition (Wang et al., 2023) for multi-hop QA, enhancing the context understanding for multi-run and conversation QA (Roy et al., 2024), facilitating the interactions across modalities for multi-modal QA (Dong et al., 2024b), and improving the explainability of the generated answers (Christmann et al., 2023). However, the remaining challenge is how to achieve efficient vector indexing and search over large-scale KGs and make a balance between the cost and performance?
@@ 4 Evaluation and Application
We summarize the metrics & dataset and showcase the related applications and demos.
@@ 4.2 Applications
We showcase the industrial and scientific applications and demos (details in Appendix §E.2) in synthesizing LLMs with KGs for QA.
@@ 5 Open Challenges and Opportunities
We summarize the challenges by highlighting the opportunities and discussing the future direction.
@@
Scaling to both Effectiveness and Efficiency. LLM+KG systems retrieve the facts and perform multi-hop reasoning under tight latency and memory budgets. Three bottlenecks are emerging: (1) Structure-aware retrieval: Vanilla dense or sparse retrieval treats a KG as an unordered triples, thereby discarding topological cues that are vital for pruning the search space (Tian et al., 2025). Hierarchical graph partitioning, dynamic neighbourhood expansion, and learned path-prior proposal networks are promising ways to expose structure to the retriever while keeping the index sub-linear. (2) Amortized reasoning: Current prompting pipelines repeatedly query the KG for every beam or CoT step. Caching subgraphs, reusing intermediate embeddings, and exploiting incremental-compute friendly hardware can mitigate the quadratic blow-up of iterative reasoning. (3) Lightweight answer validation: Most guardrails rely on large LLMs, where probabilistic logic programs, or bloom filter sketches could provide on-device verification with O(1) additional parameters. An opportunity is to design the retriever and validator that uncertainty estimates from the former guide selective execution of the latter.
@@
Knowledge Alignment and Dynamic Integration. Once a KG snapshot is injected into an LLM, it starts to become outdated. Just like real-world KGs usually involve adding new entities, deleting relations, and resolving contradictions. Future work should: (1) Quantify alignment: We lack metrics that score not only semantic overlap but also structural compatibility between parametric knowledge in the LLM and symbolic knowledge in the KG. Contrastive probing with synthetic counterfactuals or topology-aware alignment losses may fill this gap. (2) Facilitate real-time updates: Parameter-efficient tuning (e.g. LoRA modules keyed by graph deltas) and retrieval-time patching (streaming KGs with temporal indices) are early steps toward stream-time knowledge alignment. (3) Detect and resolve conflicts: Bayesian trust networks, source-aware knowledge distillation, and multi-agent debate protocols can estimate and reconcile confidence scores across modalities and sources. Incorporating these into the decoding objective is an open challenge with high pay-off.
@@
Explainable and Fairness-Aware QA. The scale of LLMs poses challenges to explainability and fairness in QA. While integrating KGs offers a path toward interpretable reasoning, it also introduces computational challenges and fairness concerns. Future work may consider the following directions: (1) Reasoning over subgraphs: Retrieving subgraphs from large-scale KGs is computationally expensive and often results in overly complex or incomprehensible explanations. Structure-aware retrieval and reranking methods should be employed to identify subgraphs consistent with the gold paths. Furthermore, CoT-based prompting can be used to guide LLMs in generating explicit reasoning steps grounded in the retrieved subgraphs. (2) Fairness-aware knowledge retrieval: LLMs can capture social biases from training data, but KGs may contain incomplete or biased knowledge. As a result, the fairness concerns remain in RAG (Wu et al., 2024b). Incorporating fairness-aware techniques into KG retrieval (e.g., reranking based on bias detection) and integrating them with counterfactual prompting can mitigate bias. (3) Multi-turn QA: Single-turn QA restricts the exploration of diverse perspectives and limits the exploration of reasoning processes. Developing multi-turn QA with retrieval strategies that can dynamically detect and adjust for bias and improve further the explainability and fairness through multi-turn interactions.

### v2

@@ 3.1.3 Aligning with Complex QA
In the approaches of synthesizing LLMs and KGs for QA, where KG plays the role of background knowledge, the retrieved relevant knowledge from the factual knowledge graph can reconcile knowledge conflicts from the multiple documents for multiple-doc QA (Wang et al., 2024d). The cross-modal reasoning can facilitate the cross-modal interaction and alignment for multi-modal QA (Suri et al., 2024). Additionally, the question decomposition of multi-hop QA can be augmented by fusing the knowledge from LLMs and KGs, which further facilitates iterative reasoning for generating the accurate final question (Saleh et al., 2024; Cao et al., 2024). The RAG (Roy et al., 2024) and KG-RAG (Sanmartin, 2024) can also improve the capabilities of LLMs in understanding the user’s interactions for generating accurate answers for conversational QA. However, the key technical challenge behind this methodology is how to retrieve the relevant knowledge from large-scale KGs and then effectively fuse with LLMs without inducing knowledge conflicts?
@@ 3.2.1 Offline KG Guidelines
In offline KG guidelines, KG supplies potential paths or subgraphs before the reasoning process by the LLM, from which the LLM selects the most relevant path for reasoning. For instance, SR (Zhang et al., 2022) trains a subgraph retriever that operates independently from the downstream reasoning process, while Keqing (Wang et al., 2023) decomposes complex questions using predefined templates, retrieves candidate entities and triples from a KG. EtD (Liu et al., 2024a) initially uses a GNN to identify promising candidates and extracts relevant fine-grained knowledge pertinent to the questions, and then creates a knowledge-enhanced multiple-choice prompt to guide LLM for generating the final answer. Recent studies have explored the application of novel formats of guidelines. GCR (Luo et al., 2024a) first converts a KG into a KG-Trie and then develops a graph-constrained decoding and a lightweight LLM to generate multiple reasoning paths and candidate answers. KELDaR (Li et al., 2024b) introduces the question decomposition and atomic retrieval modules to extract implicit information and retrieves the relevant subgraphs from KG to augment LLM for QA.
@@ 3.2.2 Online KG Guidelines
This paradigm emphasizes that the guidance of the KG directly involves the LLM’s reasoning process, rather than merely relying on existing static knowledge. For example, Oreo (Hu et al., 2022) uses a contextualized random walk across a KG and conducts a single step of reasoning through the specific layers. KBIGER (Du et al., 2022) considers the (kk-1)-th reasoning graph to construct the $k$-th instruction for reasoning and corrects erroneous predictions of intermediate entities. LLM-ARK (Huang, 2023) approaches reasoning tasks as sequential decision-making processes and employs Proximal Policy Optimization (PPO) for optimization. ToG (Sun et al., 2024a) allows LLMs to iteratively perform beam search over KGs, by which the most promising reasoning paths and the most likely reasoning outcomes are generated. In contrast, ToG-2 (Ma et al., 2025b) utilizes entities as intermediaries to guide LLMs toward precise answers based on iterative retrieval between documents and KGs. KG-CoT (Zhao et al., 2024b) leverages external KGs to generate reasoning paths for joint reasoning of LLMs and KGs to enhance the reasoning capabilities of LLMs for QA. To identify the relevant KG subgraphs, HippoRAG (Gutiérrez et al., 2024) integrates multip-hop reasoning with single-step multi-hop knowledge retrieval.
@@ 3.2.3 Agent-based KG Guidelines
This integration allows the Agent to leverage structured knowledge for augmenting the decision-making and problem-solving capabilities of LLMs. KG-Agent (Jiang et al., 2024) is a multifunctional toolbox integrating LLMs with a KG-based executor and a knowledge memory system that autonomously selects tools and updates the memory to enhance the reasoning of LLMs over KGs. ODA (Sun et al., 2024b) incorporates KG reasoning capabilities through a global observation approach, which improves reasoning abilities by employing a cyclical paradigm of observation, action, and reflection. GREASELM (Zhang et al., 2021) effectively integrates encoded representations from LMs and GNNs by introducing several modality interaction layers to seamlessly blend structured knowledge with language contexts. PoG (Chen et al., 2024a) integrates reflection and self-correction mechanisms to adaptively explore the reasoning paths over KG via an LLM agent, and then augments the LLM in complex reasoning and question answering based on the retrieved knowledge paths. ATOMR (Xin et al., 2025) leverages LLMs as a reasoning agent to retrieve and incorporate the knowledge across multiple knowledge sources for augmenting the reasoning capability of LLM in knowledge-intensive question answering.
@@ 3.2.4 Aligning with Complex QA
The approaches that incorporate KGs with LLMs can enable multi-hop and iterative reasoning over the factual KGs, further augmenting the reasoning capability of LLMs for complex QA. The challenges of knowledge retrieval and conflicts across modalities and knowledge sources, as well as complex knowledge reasoning and question answering in multi-document QA, multi-modal QA, and multi-hop QA, can be mitigated through joint reasoning over factual KGs and LLMs (Suri et al., 2024; Liang et al., 2025; Qiao et al., 2024; Xin et al., 2025). Moreover, the joint reasoning over the factual KGs and LLMs provides the logical inference chains and anchor for LLMs to generate explainable answers with clear evidence from factual KGs (Zhao et al., 2024b). Although the joint reasoning over the factual KGs and LLMs can address the challenges of complex QA well, joint KG–LLM reasoning remains inefficient because large-scale graph traversal is computationally intensive and time-consuming. Moreover, the reasoning capabilities of KGs mainly depend on the completeness and knowledge coverage of KGs, where the incomplete, inconsistent, and outdated knowledge from KGs might induce noise or conflicts. The main challenge lies in how to improve the reasoning efficiency over the large-scale graph and reasoning capabilities under incomplete KG?
@@ 3.3.1 KG-Driven Filtering and Validation
KGs enhance the accuracy and reliability of LLM outputs by filtering and validating candidate answers with structured and verified knowledge. For instance, ACT-Selection (Salnikov et al., 2023) filters and re-ranks answer candidates based on their types extracted from Wikidata. Q-KGR (Zhang et al., 2024c) improves the reasoning capabilities of LLMs by filtering out the irrelevant knowledge based on the ranking of the relevance score between the question and knowledge. KGs can improve the factual accuracy of generated answers, as demonstrated by KG-Rank (Yang et al., 2024), which integrates medical KGs with re-ranking techniques to enhance the credibility of generated responses. Moreover, KGR (Guan et al., 2024) autonomously extracts and validates factual statements, significantly boosting performance on factual QA.
@@ 3.3.2 KG-Augmented Output Refinement
Integrating KGs with LLMs is essential to refining the outputs of LLMs for greater clarity and accuracy. EFSUM (Ko et al., 2024) employs LLM as a fact summarizer to generate relevant summaries from KGs, thereby improving performance in zero-shot QA. InteractiveKBQA (Xiong et al., 2024) enables iterative knowledge interactions, allowing LLMs to generate logical forms and refine outputs based on user feedback. LPKG (Wang et al., 2024b) fine-tunes LLMs with KG-derived planning data to enhance the planning capabilities of LLMs in sophisticated reasoning for complex QA.
@@ 3.3.3 Aligning with Complex QA
The approaches that leverage the retrieved factual evidence from KGs for refinement and validation are designed to augment the capability of LLMs in understanding user interactions and verifying the intermediate reasoning for multi-hop QA (Chen et al., 2024b) and conversational QA (Xiong et al., 2024), respectively. However, the common case is that the knowledge from KGs is not available or not enough to verify the intermediate results, since the factual knowledge from KGs might be incomplete (Zhou et al., 2025). In particular, the knowledge conflicts between the intermediate answer and KG facts might induce irrelevant results due to the poorly verified intermediate results. Meanwhile, the refinement and validation of results largely depend on the correctness, timeliness, and completeness of factual knowledge in KGs. Thereby, the main challenges of this approach lie in how to handle the knowledge conflict between intermediate answers and KG facts and incrementally update KGs to ensure factual knowledge in KGs is up-to-date and correct?
@@ 3.4.1 Hybrid Method
In addition to acting as a single role, the approaches (comparison in Table 6) where KGs serve multiple roles when synthesizing with LLMs for QA have been investigated. For instance, KG-RAG (Sanmartin, 2024) introduces the Chain-of-Explorations (CoE) to rank and obtain the top-$k$ nodes or relationships based on vector similarity, by which the most relevant knowledge is selected and then fed into LLMs for the final answer. LongRAG (Zhao et al., 2024a) retrieves the top-$k$ most relevant chunks based on a hybrid retriever and analyzes their relevance with the query by introducing a CoT-guided filter. Furthermore, the KG-augmented prompting is introduced to LLMs for augmenting the generation of the final answer. In KG-Rank (Yang et al., 2024), multiple ranking methods are introduced to refine the retrieved triples for augmenting the reasoning with the most relevant knowledge. FRAG (Zhao, 2024) introduces reasoning-aware and flexible-retrieval modules to retrieve reasoning paths from KGs, thereby guiding and augmenting LLM for efficient reasoning and answer generation. KGQA (Ji et al., 2024) combines the CoT-based prompting with graph retrieval to improve the retrieval quality and the multi-hop reasoning capability of LLMs in QA.
@@ 3.4.2 Optimization
To mitigate the low efficiency and high computing costs of existing methods in synthesizing LLMs and KGs for complex QA, several optimization techniques (extension in Appendix § D) are proposed to improve the efficiency of synthesizing LLMs and KGs. (1) Index-based optimization. It aims to accelerate the process of learning embeddings and vector retrieval for given questions and knowledge context. For instance, PG-RAG (Liang et al., 2024b) proposes dynamic and adaptable knowledge retrieval indexes based on LLMs that can effectively handle complex queries and improve the overall performance of RAG systems in QA tasks. (2) Prompt-based optimization. It mainly enhances the quality of the prompts and facilitates the knowledge fusion based on prompt engineering. For example, KGP (Wang et al., 2024d) proposes a KG prompting approach to enhance the prompt for LLMs and optimize the knowledge retrieval by introducing the KG construction module and LLM-based graph traversal agent. (3) Cost-based optimization. It aims to minimize computation costs by reducing the number of calls to LLMs and accelerating knowledge retrieval. In particular, SPOKE KG-RAG (Soman et al., 2024) proposes a token-based optimized KG-RAG framework that integrates explicit and implicit knowledge from KG with LLM to enhance LLMs for cost-effective QA.
@@ 3.4.3 Aligning with Complex QA
The hybrid method that involves the multiple roles of KG when synthesizing LLMs and KGs can address the limitations and challenges of complex QA in knowledge interactions and complex reasoning across modalities and knowledge sources. The combination of knowledge fusion, RAG, CoT-based reasoning, and ranking-based refinement is capable of accelerating the complex question decomposition (Wang et al., 2023) for multi-hop QA, enhancing the context understanding for conversation QA (Roy et al., 2024), facilitating the interactions across modalities for multi-modal QA (Dong et al., 2024b), and improving the explainability of the generated answers (Christmann et al., 2023). However, the remaining challenge is how to achieve efficient vector indexing and search over large-scale KGs and maintain a balance between the cost and performance?
@@ 4 Evaluation and Application
We summarize the metrics & dataset and showcase the related applications and demonstrations.
@@ 4.2 Applications
We showcase the industrial and scientific applications and demonstrations (details in Appendix §E.2) in synthesizing LLMs with KGs for QA.
@@ 5 Open Challenges and Opportunities
Even though the synthesis of LLM and KG can take advantage of their strengths and be mutually beneficial for complex question answering, the current LLM+KG QA systems may still meet challenges. We summarize the challenges by highlighting the opportunities and discussing the future directions.
@@
Scaling to both Effectiveness and Efficiency. LLM+KG systems retrieve the factual knowledge and perform multi‐hop reasoning under tight latency and memory budgets. Three bottlenecks are emerging: (1) Structure‐aware retrieval: Vanilla dense or sparse retrieval treats a KG as an unordered triple, thereby discarding topological cues that are vital for pruning the search space (Tian et al., 2025). Hierarchical graph partitioning, dynamic neighbourhood expansion, and learned path‐prior proposal networks are promising ways to expose structure to the retriever while keeping the vector index sublinear. (2) Amortized reasoning: Current retrieval and prompting pipelines repeatedly query the KG for every Beam search or CoT step. Caching subgraphs, reusing intermediate embeddings, and exploiting incremental‐computation friendly hardware can mitigate the quadratic blow‐up of iterative reasoning. (3) Lightweight answer validation: Most guardrails rely on LLMs, while probabilistic logic programs and bloom filter sketches with KG-based fact-checking offer a lightweight solution. An opportunity is to design a retriever and validator that estimates the uncertainty of retrieval results and then guides the selective execution of validation.
@@
Knowledge Alignment and Dynamic Integration. Once a KG snapshot is injected into an LLM, it starts to become outdated, just like real‐world KGs usually involve adding new entities, deleting relations, and resolving contradictions. Future work should focus on: (1) Quantify alignment: We lack metrics that score not only semantic overlap but also structural compatibility between parametric knowledge in the LLM and symbolic knowledge in the KG. Contrastive probing with synthetic counterfactuals or topology‐aware alignment losses may fill this gap. (2) Facilitate real‐time updates: Parameter‐efficient tuning (e.g. LoRA modules keyed by graph deltas) and retrieval‐time patching (streaming KGs with temporal indices) are early steps toward stream‐time knowledge alignment. (3) Detect and resolve conflicts: Bayesian trust networks, source‐aware knowledge distillation, and multi‐agent debate protocols can estimate and reconcile confidence scores across modalities and sources. Incorporating these into the decoding objective is an open challenge with high pay‐off.
@@
Explainable and Fairness-Aware QA. The scale of LLMs poses challenges to explainability and fairness in QA. While integrating KGs offers a path toward interpretable reasoning, it also introduces computational challenges and fairness concerns. Future work may consider the following directions: (1) Reasoning over subgraphs: Retrieving subgraphs from large-scale KGs is computationally expensive and often results in overly complex or incomprehensible explanations. Structure-aware retrieval and reranking methods should be employed to identify subgraphs consistent with the gold subgraphs. Furthermore, CoT-based prompting can be used to guide LLMs in generating explicit reasoning steps grounded in the retrieved subgraphs. (2) Fairness-aware knowledge retrieval: LLMs can capture social biases from training data, but KGs may contain incomplete or biased knowledge. As a result, the fairness concerns remain in RAG (Wu et al., 2024b). Incorporating fairness-aware techniques into KG retrieval (e.g., reranking based on bias detection) and integrating them with counterfactual prompting can mitigate bias. (3) Explainable conversational QA: Single-turn QA restricts the exploration of diverse perspectives and limits the reasoning processes. Developing conversational QA with retrieval strategies can dynamically detect and adjust knowledge biases and further improve the explainability of the RAG-based QA system through multi-turn user interactions.